OpenStudy (specek18):

If f(x)=3x−6x−2, what is the average rate of change of f(x) over the interval [6, 8]?

OpenStudy (specek18):

A. 0 B. 1 C. 3 D. 3.5

OpenStudy (specek18):

those are the options, will give medals! please help!!! @abb0t @AngelWilliams16 @bradenhart @Gokuporter @Taylor<3sRin

OpenStudy (danjs):

Average rate of change is the secant line from connecting those two points

OpenStudy (specek18):

how do you find that?

OpenStudy (danjs):

find f(6) and f(8) first

OpenStudy (anonymous):

Slope of secant line

OpenStudy (danjs):

just saw that .. yeah *Slope

OpenStudy (specek18):

i'm AWFUL at math....so how would you solve for f?

OpenStudy (danjs):

\[\frac{ f(6) - f(8) }{ 6 - 8 }\]

OpenStudy (danjs):

Let X = 6 and calculate f(6 =

OpenStudy (anonymous):

Solve for \(f\)? They already gave \(f\) to you.

OpenStudy (danjs):

f(6) = 3*6 - 6*6 - 2

OpenStudy (specek18):

theres the question...i'm very confused:/

OpenStudy (danjs):

\[ave change = \frac{ f(6) - f(2) }{ 6-2 } = \frac{ -20 - (-8) }{ 6 - 2 }\]

OpenStudy (calculusfunctions):

Average rate of change of a function over the interval\[x _{1} \le x \le x _{2}\]is given by\[\frac{ f(x _{2})-f(x _{1}) }{ x _{2}-x _{1} }\]

OpenStudy (danjs):

It is just the slope of the line connecting the points at the endpoints of the interval. as above

OpenStudy (calculusfunctions):

@DanJS you're not supposed to give out the answers. You're supposed to teach.

OpenStudy (specek18):

so 3x-6/x-2....but they already gave that to me...

OpenStudy (danjs):

Which question do youw ant to do, the one you posted first , or the linked one?

OpenStudy (specek18):

it's the same question...

OpenStudy (specek18):

did i mistype something?

OpenStudy (danjs):

no, you typed it differently. ill go with the link

OpenStudy (specek18):

that's the correct one:)

OpenStudy (danjs):

so \[f(x) = \frac{ 3x - 6 }{ x - 2 }\] Evaluate that at x = 6 and at x = 8

OpenStudy (danjs):

for example f(6) = (3*6 - 6) / (6-2) = 12 / 4 = 3

OpenStudy (danjs):

now do f(8)

OpenStudy (specek18):

i got what you just did @DanJS

OpenStudy (danjs):

right , so find out what f(8) is now

OpenStudy (danjs):

Then use this \[slope = \frac{ f(6) - f(8) }{ 6 - 8 }\]

OpenStudy (danjs):

Slope is the average rate of change over that interval [6,8]

OpenStudy (jhannybean):

ok, \(f(x) = 3x−6x−2\) you want to find the average RoC, and that is found by \(\dfrac{f(b) - f(a)}{b-a}\) First find \(f(8)~,~ f(8) = 3(8)-6(8)-2\). Then find \(f(6) ~,~ f(6)= 3(6)-6(8)-2\) Then just plug it into your formula.

OpenStudy (jhannybean):

@DanJS , you're solving it like \(\dfrac{f(a)-f(b)}{a-b}\), seems a little opposite.

OpenStudy (danjs):

it doesnt matter

OpenStudy (jhannybean):

Oh, hm, thought it would make a difference in the slope.

OpenStudy (danjs):

nah

OpenStudy (danjs):

neg/neg = pos/pos, either way you look at it

OpenStudy (danjs):

So where you at with it SPECEK

OpenStudy (specek18):

sorry trying to keep up with everyone!:) so 3 would be the average rate of change?

OpenStudy (danjs):

not sure, i didnt calculate it, what did you do

OpenStudy (specek18):

im going to post a pic. one sec:)

OpenStudy (danjs):

ok, ill type the result in the mean time

OpenStudy (specek18):

that's all i've got..

OpenStudy (danjs):

\[f(8) = 0\] f(6) = 0 \[\frac{ f(8) - f(6) }{ 8 - 6 } = \frac{ 3 - 3 }{ 2 } = \frac{ 0 }{ 2 } = 0\]

OpenStudy (danjs):

It is a horizontal line actually, the rate of change is zero

OpenStudy (danjs):

f(8) = 3 f(6) = 3

OpenStudy (specek18):

huh....i do understand how you did that:) thank you so much! could you help me with a few more problems?

OpenStudy (danjs):

i put zero in the one above on accident for f(8) and f(3) they are both 3

OpenStudy (danjs):

Here look at this, it can help you

OpenStudy (specek18):

wait so the rate of change is 3 or 0???

OpenStudy (danjs):

\[f(x) = \frac{ 3x - 6 }{ x - 2 } = \frac{ 3(x-2) }{ (x-2) } = 3\]

OpenStudy (danjs):

It is a horizontal line at y = 3 ...f(x) = 3

OpenStudy (specek18):

okay so 3? want to make sure!!

OpenStudy (danjs):

no the rate of change of a horizontal line, is zero

OpenStudy (specek18):

alright. sorry all of this confuses me! i'm going with 0!

OpenStudy (danjs):

the y coordinate does not change as the x coordinate changes in a horizontal line, the slope is zero

OpenStudy (danjs):

here to summarize...

OpenStudy (danjs):

f(x) = (3x-6)/(x-2) Average rate of change btween two points is the slope. \[slope = \frac{ change of y }{ change of x } = \frac{ f(8) - f(6) }{ 6-8 } = \frac{ 3 - 3 }{ 2 } = \frac{ 0 }{ 2} = 0\] The average rate of change of f(x) over the interval [6,8] is ZERO.

OpenStudy (danjs):

That is it

OpenStudy (specek18):

okay:) thank you!!

OpenStudy (specek18):

heres my next question...

OpenStudy (specek18):

@DanJS

OpenStudy (danjs):

This would be the same thing... \[ave change = \frac{ f(3) - f(-2) }{ 3 - (-2) }\]

OpenStudy (danjs):

They gave you a table of values, so you dont have to calculate f(3) and f(-2)

OpenStudy (specek18):

i have trouble setting it up. lemme confirm the answer with you?:)

OpenStudy (danjs):

type out the first step

OpenStudy (specek18):

plugging in the values?

OpenStudy (danjs):

the interval is [-2 , 3] find f(-2) and f(3) from the table

OpenStudy (specek18):

i got 2.5/5

OpenStudy (specek18):

which equals half, .5

OpenStudy (specek18):

correct?

OpenStudy (danjs):

no

OpenStudy (danjs):

watch..

OpenStudy (danjs):

The interval is over x=-2 to x=3 , [-2, 3] from table f(-2) = 2.5 f(3) = 5 average Rate of Change over [-2,3] = \[\frac{ f(3) - f(-2) }{ 3 - (-2) } = \frac{ 5 - 2.5 }{ 3 + 2 }\]

OpenStudy (danjs):

oh i see, you just gave me the final answer... yeah it is 2.5/5

OpenStudy (danjs):

I thought you just divided the f(-2) and f(3)

OpenStudy (specek18):

yeah i'm sorry! i have trouble explaining myself.

OpenStudy (specek18):

but the final would be half?

OpenStudy (danjs):

So all you need to remember for the AVERAGE RATE OF CHANGE y = f(x); over an interval [a,b] Ave change = \[\frac{ f(b) - f(a) }{ b - a }\] average change is found by evaluating the function f(x) at x=b and x=a, then using the above formula.

OpenStudy (danjs):

yeah final 2.5/5 = 1/2

OpenStudy (specek18):

okay...thank you so much for your patience with me. it means SOO much! @DanJS i just have two more questions...

OpenStudy (specek18):

of course! thank you lots